ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 110, 1999

Berkovich, Y. Groups with few characters of small degrees  325

Caranti, A. Loop algebras of Zassenhaus algebras in characteristic three

61

Chaber, J. and R. Pol On the Cantor–Bendixson derivative, resolvable ranks, and perfect set theorems of A. H. Stone

103

Chen, B.-Y. and L. Vrancken CR-submanifolds of complex hyperbolic spaces satisfying a basic equality

341

Göbel, R. and D. Simson Rigid families and endomorphism algebras of Kronecker modules

293

Koldobsky, A. A generalization of the Busemann–Petty problem on sections of convex bodies

75

Loeb, J. J. and M. Nicolau On the complex geometry of a class of non-Kählerian manifolds

371

Longobardi, P., M. Maj and A. Mann Minimal classes and maximal class in p-groups

93

Maj, M. See P. Longobardi

Mann, A. See P. Longobardi

Nicolau, M. See J. J. Loeb

Nies, A. and A. Sorbi Branching in the enumeration degrees of the

Σ_{2}^{0} sets 29

Oh, H. On discrete subgroups containing a lattice in a horospherical subgroup

333

Pol, R. See J. Chaber

Puczyłowski, E. R. and A. Smoktunowicz The nil radical of power series rings

317

Shah, R. The central limit problem on locally compact groups  189

Simson, D. See R. Göbel

Smoktunowicz, A. See E. R. Puczyłowski

Sorbi, A. See A. Nies

Tanaka, N. Global solutions of abstract quasi-linear evolution equations of “hyperbolic” type

219

Thomsen, K. The variational principle for the defect of factor maps 359

Visintin, B. Carathéodory balls and norm balls in a class of convex bounded Reinhardt domains

1

Volpert, A. I. See V. A. Volpert

Volpert, V. A. and A. I. Volpert Existence and stability of multidimensional travelling waves in the monostable case

269

Vrancken, L. See B.-Y. Chen

Wolf, R. Averaging distances in real quasihypermetric Banach spaces of finite dimension

125

Young, L.-S. Recurrence times and rates of mixing  153

Zippin, M. Almost locally minimal projections in finite dimensional Banach spaces 253